### v2

# Semantic Energy: Detecting LLM Hallucination Beyond Entropy

Abstract
Large Language Models (LLMs) are being increasingly deployed in real-world applications, but they remain susceptible to hallucinations, which produce fluent yet incorrect responses and lead to erroneous decision-making. Uncertainty estimation is a feasible approach to detect such hallucinations. For example, semantic entropy estimates uncertainty by considering the semantic diversity across multiple sampled responses, thus identifying hallucinations. However, semantic entropy relies on post-softmax probabilities and fails to capture the model’s inherent uncertainty, causing it to be ineffective in certain scenarios. To address this issue, we introduce Semantic Energy, a novel uncertainty estimation framework that leverages the inherent confidence of LLMs by operating directly on logits of penultimate layer. By combining semantic clustering with a Boltzmann-inspired energy distribution, our method better captures uncertainty in cases where semantic entropy fails. Experiments across multiple benchmarks show that Semantic Energy significantly improves hallucination detection and uncertainty estimation, offering more reliable signals for downstream applications such as hallucination detection. The code and intermediate data are available at https://github.com/SemanticEnergy.
1 Introduction
Large Language Models (LLMs) have been widely deployed in various aspects of production and daily life, demonstrating strong capabilities in different fields (Schlegel et al., 2025; Xiang et al., 2025). However, LLMs are still prone to being influenced by hallucinations and are prone to generate incorrect answers in situations where they lack knowledge, thus misleading users into making errors (Zhou et al., 2024; Farquhar et al., 2024). Recently, uncertainty estimation has been shown to be a reliable indicator for detecting hallucinations, reflecting the tendency of an LLM to generate hallucinations (Xiao & Wang, 2021; Huang et al., 2024). When the uncertainty of an LLM response is high, it often suggests a greater likelihood that the response is a hallucination, prompting further actions such as self-reflection (Renze & Guven, 2024; Kirchhof et al., 2025), regenerating of answers (Xu et al., 2025), or intervention by human experts (Liu et al., 2025; Hopkins et al., 2025).
Entropy is a commonly used metric for estimating uncertainty in LLM (Cheng et al., 2025; Duan et al., 2024). Similarly to traditional discriminative models, high entropy indicates high uncertainty because it means that the model cannot confidently select a particular outcome. However, due to the nature of natural language, the entropy of a single response cannot accurately reflect the reliability of LLMs. Specifically, even though LLMs may not confidently generate the next token, the semantic meaning of any generated token can still be the same. In such cases, we cannot identify an unreliable response simply attributing to its low probability of being generated. To accurately describe the uncertainty of responses composed of natural language, semantics must be considered.
Semantic entropy (Farquhar et al., 2024) is a typical method to characterize the semantic uncertainty of responses, effectively representing the probability that an LLM generates hallucinations. Given a question, semantic entropy involves sampling multiple responses, clustering them based on their semantic meaning, and then replacing individual responses with clusters to calculate entropy, thus achieving semantic-aware uncertainty characterization. Based on this method, a wide range of downstream applications have been developed, such as guiding Chain-of-Thought (CoT) reasoning (Ye et al., 2025) and parallel thinking (Xu et al., 2025). However, semantic entropy has significant drawbacks stemming from entropy itself: it fails to capture the model’s inherent uncertainty, leading to its ineffectiveness in some scenarios.
A representative case occurs when the model produces identical responses in multiple sampling instances for a given question, as illustrated in Fig. 1. According to semantic entropy, the resulting value is 0, which is considered a reliable response. However, even answering incorrectly, LLMs might also provide responses with the same semantics. Among samples with consistently semantic responses across multiple responses, the proportion of incorrect responses (like Question3 in Fig. 1) approaches 50% in some datasets. In such cases, it is necessary to leverage the model’s inherent uncertainty for differentiation: even if the LLM provides multiple responses with the same semantics for two different questions, their corresponding reliability still differs. In scenarios with a higher inherent uncertainty in the model, the likelihood of the LLM making mistakes is greater.
Several previous studies have shown that logits exhibit stronger inherent capabilities to characterize uncertainty compared to probabilities, and the magnitude of logits can indicate whether the model has undergone adequate training in a given scenario (Liu et al., 2020; Fu et al., 2025; Zhang et al., 2024). For example, in out-of-distribution (OOD) detection, studies have highlighted that the logit values for in-distribution (InD) samples are significantly higher than those for OOD samples (Liu et al., 2020). Recent work named LogToKU (Ma et al., 2025) points out that probabilities lose the intensity information of logits during normalization, thus limiting their ability to represent the inherent uncertainty of LLM. From this insight, we propose a new method to improve the failure cases of Semantic Entropy, termed Semantic Energy. Specifically, for a given prompt, we first perform multiple response samplings, followed by semantic sampling. When calculating the final uncertainty, rather than relying on probability as in Semantic Entropy, we estimate the response uncertainty based on logits, enabling the estimated uncertainty to reflect the model’s inherent uncertainty. Our proposed metric significantly outperforms Semantic Entropy in evaluating the reliability of LLM responses, particularly in scenarios where Semantic Entropy fails. The main contributions are as follows:
  • We expose the limitations of current uncertainty estimation methods based on probability and identify the failure cases in Semantic Entropy.
  • We introduce Semantic Energy, a novel framework to evaluate the uncertainty of LLM responses, which indicates potential errors in the responses.
  • We instantiate Semantic Energy using the Boltzmann formulation, and in the hallucination detection task, it achieves an average performance improvement of more than 13% compared in terms of AUROC to Semantic Entropy in cases where the latter is confident.
2 Preliminaries
Figure 1: An intuitive comparison between Semantic Entropy and Semantic Energy in their ability to characterize uncertainty. Both approaches first sample and perform semantic clustering over distinct clusters of activity. The difference lies in the computation: Semantic Entropy is calculated based on normalized probabilities, while Semantic Energy is derived from logits. This enables Semantic Energy to distinguish cases when Semantic Entropy fails.
2.1 Estimating LLM Uncertainty with Token-Level Entropy
Let $\bm{q}$} denote a natural language query provided as input to the LLM. Given the prompt $\bm{q}$}, a single response is generated as:
${\bm{x}}=[x_{1},x_{2},\dots,x_{T}],$ (1)
where $\bm{x}$} is a token sequence of length $T$. The model assigns a conditional probability to each token $x_{t}$ given the previous context and the input query $p(x_{t}\mid x_{<t},{\bm{q}})$, and the token-level entropy for position $t$ is defined as:
$H_{t}=-\sum_{x\in\mathcal{V}}p(x\mid x_{<t},{\bm{q}})\log p(x\mid x_{<t},{\bm{q}}),$ (2)
where $\mathcal{V}$ is the model vocabulary. A common approach to estimating the uncertainty of the full response is to aggregate the token-level entropy. This can be estimated by computing the average:
$H_{\mathrm{avg}}({\bm{x}})=\frac{1}{T}\sum_{t=1}^{T}H_{t},$ (3)
where $H_{\mathrm{avg}}({\bm{x}})$ represents the uncertainty of the response, with a high entropy indicating greater uncertainty. Considering that different tokens have varying levels of importance for this response, some works (Duan et al., 2024) consider using a weighted average to increase the weight of critical tokens when calculating the entropy:
$H_{\mathrm{wavg}}({\bm{x}})=\sum_{t=1}^{T}w_{t}H_{t},\quad\text{where }\sum_{t=1}^{T}w_{t}=1.$ (4)
Although token-level entropy provides a fine-grained view of local uncertainty during generation, it may not fully capture the diversity in the meanings of different responses. This motivates a higher-level notion of uncertainty based on the distribution over semantically distinct outputs, leading to the concept of semantic entropy (Kuhn et al., 2024).
2.2 Semantic Entropy and Response Clustering
To capture semantic-level uncertainty, semantic entropy samples a set of $n$ candidate responses to the query $\bm{q}$} from LLM:
${\mathbb{X}}=\{{\bm{x}}^{(1)},{\bm{x}}^{(2)},\dots,{\bm{x}}^{(n)}\},\quad{\bm{x}}^{(i)}=[x^{(i)}_{1},x^{(i)}_{2},\dots,x^{(i)}_{T_{i}}$], (5)
where ${\bm{x}}^{(i)}$ indicates the $i$-th sampled response and $T_{i}$ indicates the number of tokens in this response. Each response ${\bm{x}}^{(i)}$ has an associated likelihood in the model:
$p({\bm{x}}^{(i)}\mid{\bm{q}})=\prod_{t=1}^{T_{i}}p(x_{t}^{(i)}\mid x_{<t}^{(i)},{\bm{q}}),\quad\bar{p}({\bm{x}}^{(i)})=\frac{p({\bm{x}}^{(i)}\mid{\bm{q}})}{\sum_{j=1}^{n}p({\bm{x}}^{(j)}\mid{\bm{q}})}.$ (6)
where $\bar{p}({\bm{x}}^{(i)})$ indicates a normalized distribution over the sampled responses. Due to surface-level variability in language, semantically similar responses may have different forms. Therefore, semantic entropy clusters the responses into $K$ semantically coherent groups:
${\mathbb{C}}=\{{\mathbb{C}}_{1},{\mathbb{C}}_{2},\dots,{\mathbb{C}}_{K}\},\quad{\mathbb{C}}_{k}\subseteq{\mathbb{X}},$ (7)
where each cluster ${\mathbb{C}}_{k}$ contains responses that are semantically equivalent. The probability mass assigned to each cluster is defined as the sum over its members:
${p}({\mathbb{C}}_{k})=\sum_{{\bm{x}}^{(i)}\in{\mathbb{C}}_{k}}\bar{p}({\bm{x}}^{(i)}).$ (8)
The semantic entropy is then given by:
$H_{\mathrm{SE}}=-\sum_{k=1}^{K}{p}({\mathbb{C}}_{k})\log{p}({\mathbb{C}}_{k}),$ (9)
which quantifies the model’s uncertainty over distinct meanings conveyed by its responses. However, semantic entropy fails in some scenarios due to the limitations of entropy-based uncertainty estimation.
3 Modeling Uncertainty via Semantic Energy
3.1 Limitations of Entropy-Based Uncertainty Estimation
While $H_{\mathrm{SE}}$ captures semantic variability, it only reflects aleatoric uncertainty—uncertainty arising from intrinsic randomness in the generation process. However, it fails to capture epistemic uncertainty—uncertainty stemming from the model’s lack of knowledge. For example, as illustrated by the pair of instances in Low uncertainty cases in Semantic Entropy (see Fig. 1):
  • (1)
    Consider two queries ${\bm{q}}_{2}$ and ${\bm{q}}_{3}$, where the model has been extensively trained on data related to ${\bm{q}}_{2}$ (thus confident), but has limited exposure to ${\bm{q}}_{3}$ (thus uncertain).
  • (2)
    Assume that each query is sampled to obtain $5$ responses, which were subsequently grouped into a certain cluster based on their semantic similarity ($K=1$), respectively.
  • (3)
    In this case, $H_{\mathrm{SE}}=0$ for both, despite the LLM outputs $5$ incorrect answers on ${\bm{q}}_{3}$ with the same semantic.
This means that if an LLM gives a wrong answer, and the semantics of multiple sampled results are all aligned with that wrong answer, then semantic entropy will mistakenly identify it as reliable. Unfortunately, LLMs are very good at “steadfastly” repeating the same wrong response. In some datasets, nearly half of the samples with repeated identical semantics are actually incorrect answers, making this limitation impossible to ignore.
The reason behind this is that the entropy calculated based on probabilities only reflects the relative likelihood of a particular LLM response compared to other possible responses generated by the model, rather than the actual probability of that response to the question in the real world. However, when computing the probability of the next token, the model approximates the partition function as the sum of probabilities over its vocabulary. Therefore, for entropy to accurately represent uncertainty, two assumptions must hold: (1) the model has seen all possible responses (that is, the training distribution matches the real-world distribution perfectly), and (2) the model has fit the training distribution without bias (that is, the model output distribution matches the training distribution exactly). Clearly, neither of these assumptions holds. Worse still, the LLM response is based on the joint probability prediction of many tokens, rather than the single word classification of traditional discriminative models, which causes the error of the approximated partition function to accumulate to a degree that can no longer be ignored.
3.2 Energy-Based Confidence Estimation
To address this limitation, we introduce an energy-based formulation that complements semantic entropy and captures epistemic uncertainty. In thermodynamics of physics, lower energy corresponds to a more stable and less random state. Drawing on thermodynamic analogies, we treat lower-energy states as higher-confidence predictions, following the intuition that physical systems evolve toward minimal-energy configurations.
3.2.1 Boltzmann Distribution
The classical Boltzmann distribution defines the probability that a system occupying a state is capable of generating $x_{t}^{(i)}$ as:
$p(x_{t}^{(i)})=\frac{e^{-E_{t}^{(i)}}/k\tau}{Z_{t}},$ (10)
where $k$ is the Boltzmann constant, $T$ is the temperature, and $E_{t}^{(i)}$ is the token energy $x_{t}^{(i)}$, and $Z_{t}$ is the partition function. Specifically, for LLMs, $Z_{t}=\sum_{x\in\mathbb{V}}e^{-E_{t}(x)}$ is the normalization value across the entire vocabulary $\mathbb{V}$ (the difference between $\mathcal{V}$ and $\mathbb{V}$ is that $\mathcal{V}$ is the vocabulary in the predefined tokenizer of a specific LLM, while $\mathbb{V}$ is the space of possible next tokens in the real world, which is infinite and intractable). For simplicity, we assume that $Z_{t}$ is constant across $t$. The probability of a complete sequence and the average sequence-level energy can be represented as:
$p({\bm{x}}^{(i)})=\prod_{t=1}^{T_{i}}p(x_{t}^{(i)})=\frac{e^{-\sum_{t=1}^{T_{i}}E_{t}^{(i)}}}{\prod_{t=1}^{T_{i}}Z_{t}},\quad E({\bm{x}}^{(i)})=\frac{1}{T_{i}}\sum_{t=1}^{T_{i}}E_{t}^{(i)}.$ (11)
Suppose that we want to evaluate the total energy of a set $\mathbb{C}$}. According to the Boltzmann equation, the total energy of $\mathbb{C}$} is the sum of its different states:
$E_{\mathrm{Bolt}}({\mathbb{C}})=\sum_{{\bm{x}}^{(i)}\in{\mathbb{C}}}E({\bm{x}}^{(i)}).$ (12)
Lower energy indicates that the cluster containing this response is more stable, i.e., it has lower uncertainty and thus higher reliability.
3.2.2 Specific Implementation in LLMs
For a model parameterized by $\bm{\theta}$ with a vocabulary $\mathcal{V}$, the token-level energy distribution becomes:
$p(x_{t}^{(i)},\bm{\theta})=\frac{e^{-E(x_{t}^{(i)},\bm{\theta})/k\tau}}{Z_{\bm{\theta}}},\quad Z_{\bm{\theta}}=\sum_{x_{t}^{(i)}\in\mathcal{V}}e^{-E(x_{t}^{(i)},\bm{\theta})/k\tau},$ (13)
where $Z_{\bm{\theta}}$ indicates the partial function of the model parameterized by $\bm{\theta}$. Under a Bayesian treatment, total uncertainty is obtained by marginalizing over $\bm{\theta}$:
$p(x_{t}^{(i)},\mathcal{D})=\int_{\bm{\theta}}p(x_{t}^{(i)},\bm{\theta})\cdot{\tt p}(\bm{\theta}\mid\mathcal{D})\,d\bm{\theta}.$ (14)
However, we cannot obtain all possible $\bm{\theta}$, but we can only estimate uncertainty under specific models. For any given $\bm{\theta}$, the conditional probability of generating a token from a subset vocabulary $\mathcal{V}\subseteq\mathbb{V}$ is:
$p(Z_{\bm{\theta},t}\mid Z_{t})=\frac{\sum_{x_{t}^{(i)}\in\mathcal{V}}e^{-E(x_{t}^{(i)},\bm{\theta})/k\tau}}{\sum_{x_{t}^{(i)}\in\mathbb{V}}e^{-E(x_{t}^{(i)},\bm{\theta})/k\tau}}.$ (15)
Therefore, given the model parameterized by $\bm{\theta}$, Eq. 14 can be approximated as:
$\tilde{p}(x_{t}^{(i)},\mathcal{D})=p(x_{t}^{(i)},\theta)\cdot p(Z_{\bm{\theta},t}\mid Z_{t})=\frac{e^{-E(x_{t}^{(i)},\bm{\theta})/k\tau}}{Z_{t}},$ (16)
where $\tilde{p}(x_{t}^{(i)},\mathcal{D})$ indicates a sampled approximation of $p(x_{t}^{(i)},\mathcal{D})$. Similarly to softmax-based probability modeling, the probability of a complete sequence is the joint probability of all tokens in the entire sequence:
$\tilde{p}({\bm{x}}^{(i)})=\prod_{t=1}^{T_{i}}\tilde{p}(x_{t}^{(i)})=\frac{e^{-\sum_{t=1}^{T_{i}}\tilde{E}_{t}^{(i)}/k\tau}}{\prod_{t=1}^{T_{i}}Z_{t}},\quad\tilde{E}({\bm{x}}^{(i)})=\frac{1}{T_{i}}\sum_{t=1}^{T_{i}}\tilde{E}_{t}^{(i)},$ (17)
where $\tilde{E}({\bm{x}}^{(i)})$ indicates the average sequence-level energy, which can also be interpreted as the energy per unit volume in thermodynamics, that is, the free energy density (Callen, 1993). To extend to semantic clusters, we treat each cluster ${\mathbb{C}}_{k}$ energy as a scaled joint energy:
$\tilde{E}_{\mathrm{Bolt}}({\mathbb{C}}_{k})=\frac{1}{n}\sum_{{\bm{x}}^{(i)}\in{\mathbb{C}}_{k}}\tilde{E}({\bm{x}}^{(i)}),$ (18)
where $n$ are the sampling times for every question.
For an LLM trained with cross-entropy loss, we represent $E(x_{t}^{(i)},\bm{\theta})$ as the negative value of the logit, that is, $E(x_{t}^{(i)},\bm{\theta})=-z_{\bm{\theta}}(x_{t}^{(i)})$, and $k\tau$ is by default the temperature used during LLM training, that is, $k\tau=1$. The final uncertainty is defined as:
$U({\bm{x}}^{(i)})=\frac{1}{nT_{i}}\sum_{{\bm{x}}^{(i)}\in{\mathbb{C}}_{k}}\sum_{t=1}^{T_{i}}-z_{\bm{\theta}}(x_{t}^{(i)}),$ (19)
which means that the lower the energy, the lower the uncertainty.
4 Experiments
4.1 Setup
Model & Baseline.
We conduct experiments using the Qwen3‑8B model (Yang et al., 2025), and the ERNIE-21B-A3B model (MOE architecture) (Baidu, 2025). Our primary goal is to highlight the differences between probability-based methods and energy-based approaches. Therefore, we use the semantic entropy (Farquhar et al., 2024) as a baseline.
Datasets & Metrics.
Experiments are performed on standard open-domain QA datasets in both Chinese and English: the Chinese dataset CSQA (He et al., 2024) and the English dataset TriviaQA (Joshi et al., 2017). To assess whether the estimated uncertainty can capture the risk that the model makes errors, we estimate the AUROC between uncertainty scores and correctness (i.e., whether the answer is correct).
4.2 Main Results
Table 1: Uncertainty estimation performance on OpenQA Datasets.
Model Dataset Semantic Entropy Semantic Energy
AUROC AUPR FPR95 AUROC(↑) AUPR(↑) FPR95(↓)
Qwen3-8B CSQA $71.6\%$ $53.6\%$ $77.0\%$ $76.1\%$ $\Uparrow_{4.5\%}$ $61.4\%$ $\Uparrow_{7.8\%}$ $74.6\%$ $\Uparrow_{2.4\%}$
TriviaQA $69.6\%$ $73.5\%$ $79.1\%$ $74.8\%$ $\Uparrow_{5.2\%}$ $79.2\%$ $\Uparrow_{5.7\%}$ $74.7\%$ $\Uparrow_{4.4\%}$
ERNIE-21B-A3B CSQA $77.4\%$ $73.2\%$ $70.9\%$ $80.2\%$ $\Uparrow_{2.8\%}$ $77.5\%$ $\Uparrow_{4.3\%}$ $65.0\%$ $\Uparrow_{5.9\%}$
TriviaQA $75.1\%$ $85.0\%$ $69.9\%$ $81.0\%$ $\Uparrow_{5.9\%}$ $89.9\%$ $\Uparrow_{4.9\%}$ $63.7\%$ $\Uparrow_{6.2\%}$
Table 1 summarizes the performance of uncertainty estimation methods on the CSQA and TriviaQA datasets. We evaluate models using standard metrics: AUROC, AUPR, and FPR@95, based on whether the uncertainty score can discriminate correct from incorrect responses.
In both models and datasets, semantic energy consistently outperforms semantic entropy. On CSQA, the Boltzmann energy improves AUROC from $71.6\%$} to $76.1\%$} on Qwen3-8B and from $77.4\%$} to $80.2\%$} on ERNIE-21B-A3B. Similar trends are observed on TriviaQA, where Boltzmann energy yields AUROC gains of more than $5\%$ compared to the semantic entropy. Improvements are also reflected in AUPR and FPR@95, indicating better calibration and reduced false positive rates.
These results highlight the robustness of energy-based uncertainty estimation, particularly in low-diversity scenarios where entropy becomes degenerate (details in Table 2). By incorporating internal model states via logits, semantic energy captures a richer signal for uncertainty estimation beyond probability-based entropy.
4.3 Ablation Studies
4.3.1 Results on questions with single cluster
Table 2: Uncertainty estimation performance on questions with single cluster.
Model Dataset Semantic Entropy Semantic Energy
AUROC AUPR FPR95 AUROC(↑) AUPR(↑) FPR95(↓)
Qwen3-8B CSQA $0.0\% 55$$5.8\%$ $95.0\%$ $66.7\%$ $\Uparrow_{16.7\%}$ $67.6\%$ $\Uparrow_{11.8\%}$ $80.3\%$ $\Uparrow_{14.7\%}$
TriviaQA $50.0\%$ $75.1\%$ $95.0\%$ $62.1\%$ $\Uparrow_{12.1\%}$ $81.6\%$ $\Uparrow_{6.5\%}$ $86.9\%$ $\Uparrow_{8.1\%}$
ERNIE-21B-A3B CSQA $50.0\%$ $77.0\%$ $95.0\%$ $58.9\%$ $\Uparrow_{8.9\%}$ $81.9\%$ $\Uparrow_{4.9\%}$ $88.4\%$ $\Uparrow_{6.6\%}$
TriviaQA $50.0\%$ $85.9\%$ $95.0\%$ $65.8\%$ $\Uparrow_{15.8\%}$ $91.4\%$ $\Uparrow_{5.5\%}$ $83.4\%$ $\Uparrow_{11.6\%}$
In Table 2, we present the case where all responses share the same semantics, that is, all responses are clustered into a single group as described in Sec. 3.1. In this scenario, semantic entropy completely fails, whereas semantic energy is still able to provide a certain level of distinction, resulting in semantic energy achieving an average performance improvement of more than 13% compared in terms of AUROC to semantic entropy in cases where the latter is confident. It is important to note that the value of semantic entropy in such cases is always zero, meaning that its performance reflects the expected behavior when the uncertainty indicator is meaningless, for example, the AUPR corresponds to the number of positive samples (i.e. correct responses).
4.3.2 Results on think Mode
We conduct experiments on Qwen-8B using the CSQA dataset and explore the case where the think mode is enabled. Specifically, we activate the think mode but discard the content within <think>...</think> during evaluation, considering only the response portion. The final results, shown in Fig. 2, are consistent with the observations in Table 1.
(a) Semantic Entropy
(b) Semantic Energy
Figure 2: Comparison of semantic entropy vs. semantic energy on CSQA datasets with think mode on.
4.3.3 Advantages of semantic cluster
Inspired by semantic entropy, we incorporate semantic influence when computing energy. If semantics are not considered and the energy of a single response is directly used to characterize the reliability of an LLM’s reply, such as in LogTokU (Ma et al., 2025), a clear problem arises: a single response having high energy does not necessarily mean that the entire cluster of responses sharing the same semantics also has high energy. This is because different responses belonging to the same semantic cluster can still have varying energy values. Therefore, we represent the energy of a response by the energy of the cluster to which it belongs. As shown in Fig. 3, we conduct an ablation study on whether to include semantics. The experimental results demonstrate that incorporating semantics significantly improves the accuracy of uncertainty estimation.
(a) TriviaQA (w/o semantic)
(b) CSimpleQA (w/o semantic)
(c) TriviaQA (with semantic)
(d) CSimpleQA (with semantic)
Figure 3: Comparison of semantic vs. non-semantic uncertainty modeling on TriviaQA and CSimpleQA datasets.
5 Related Work
Uncertainty estimation methods.
Recently, numerous uncertainty estimation methods for LLMs have been proposed. These include methods that utilize natural language for uncertainty feedback, including heuristically designed and trained approaches (Tao et al., 2025; Xiong et al., 2023; Lin et al., 2023); methods that estimate uncertainty based on model states, including those leveraging prior knowledge or statistical observations of model states (Kostenok et al., 2023; Li et al., 2025; Liu et al., 2024), or observing changes under perturbations (Zhang et al., 2025b; Gao et al., 2024); and methods that take into account the semantics of the response, including consistency-based uncertainty characterizations (Lyu et al., 2025; Bartsch et al., 2023; Xiao et al., 2025) and approaches that integrate semantics with model states (Kuhn et al., 2024; Grewal et al., 2024).
Uncertainty-guided applications.
The utilization of uncertainty estimation is widely applied in both the post-train and inference phases of LLMs. For example, minimizing entropy during the reinforcement learning process helps reduce uncertainty (Zhang et al., 2025a; Agarwal et al., 2025) and encourages exploration of critical positions with higher uncertainty (Zheng et al., 2025; Cheng et al., 2025). In the inference phase, uncertainty can guide the model’s reasoning path (Ye et al., 2025; Wang et al., 2022), guiding RAG (Guo et al., 2025; Chen & Varoquaux, 2025), collaboration (Dey et al., 2025; Kruse et al., 2025; Cui et al., 2025), and indicate when the model should stop or skip thinking (Xu et al., 2025; Zhu et al., 2025).
6 Discussion and Conclusion
In this paper, we introduce the concept of semantic energy as an enhancement to semantic entropy, an uncertainty modeling method that substitutes entropy with energy (derived from logits). Semantic energy can effectively compensate for the shortcomings of semantic entropy and better depict the inherent uncertainty within models. We elucidate the limitations imposed by probability normalization and demonstrate the potential to overcome these constraints. However, it is important to note that semantic energy is not a flawless, ultimate solution. Given that the cross-entropy loss employed in current LLM training is invariant to the scale of the logits, logits are not strictly equivalent to energy. Instead, logits exhibit energy-like characteristics only due to implicit constraints arising from network initialization and regularization during training. Therefore, to enable models to more precisely capture their own uncertainty and acknowledge their unknowns, it may be necessary to tackle the limitations introduced by cross-entropy loss during the training process.